OpenStudy (anonymous):

Second opinion needed , wanted to make sure it's correct before I submit exam.

OpenStudy (anonymous):

I got x = 0, solution is extraneous

OpenStudy (anonymous):

@austinL

OpenStudy (austinl):

\(-4\sqrt{x-3}=12\) \(\sqrt{x-3}=-3\) \(x-3=9\) \(\color{blue}{x=12}\)